# Anomaly Detection and RFI Classification with Unsupervised Learning in Narrowband Radio Technosignature Searches

### v1

@@ Abstract
The search for radio technosignatures is an anomaly detection problem: candidate signals represent needles of interest in the proverbial haystack of radio-frequency interference (RFI). Current search frameworks find an enormity of false-positive signals, especially in large surveys, requiring manual follow-up to a sometimes prohibitive degree. Unsupervised learning provides an algorithmic way to winnow the most anomalous signals from the chaff, as well as group together RFI signals that bear morphological similarities. We present GLOBULAR (Grouping Low-frequency Observations By Unsupervised Learning After Reduction) clustering, a signal processing method that uses HDBSCAN to reduce the false-positive rate and isolate outlier signals for further analysis. When combined with a standard narrowband signal detection and spatial filtering pipeline, such as turboSETI, GLOBULAR clustering offers significant improvements in the false-positive rate over the standard pipeline alone, suggesting dramatic potential for the amelioration of manual follow-up requirements for future large surveys. By removing RFI signals in regions of high spectral occupancy, GLOBULAR clustering may also enable the detection of signals missed by the standard pipeline. We benchmark our method against the Choza et al. (2024) turboSETI-only search of 97 nearby galaxies at $L$-band, demonstrating a false-positive hit reduction rate of 93.1% and a false-positive event reduction rate of 99.3%.
@@
software: NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), SciPy (Virtanen et al., 2020), Scikit-Learn (Pedregosa et al., 2011), HDBSCAN (McInnes et al., 2017), turboSETI (Enriquez & Price, 2019), setigen (Brzycki et al., 2022), SHAP (Lundberg & Lee, 2017)
@@ 1 Introduction
Since the work of Cocconi & Morrison (1959) and Drake (1961), radio frequencies have comprised the most widely explored domain in the search for extraterrestrial intelligence (SETI) due to their favorably low extinction across cosmic distances and the known efficacy of their use for communication by at least one spacefaring civilization—humanity. Since 2015, the Breakthrough Listen (BL) Initiative (Worden et al., 2017) has used various facilities, including the Robert C. Byrd Green Bank Telescope (GBT), to conduct radio-frequency searches of numerous targets, ranging in scale from the planetary (e.g., Traas et al., 2021; Franz et al., 2022) to the galactic (e.g., Gajjar et al., 2021; Choza et al., 2024), for unambiguously artificial signals, or “technosignatures.” Such signals are generally assumed to be narrowband, with signal widths on the order of 1–10 Hz, for transmission efficiency and since no known astrophysical process naturally produces emission so narrowly peaked. A high duty cycle is also typically assumed for feasibility of follow-up observations and while calculating transmitter distributions.
@@
Past GBT searches (e.g., Enriquez et al., 2017; Price et al., 2020) have primarily used two criteria to reject RFI. The first is that an extraterrestrial narrowband signal must exhibit a drift in frequency over time due to the relative motion between the transmitter and the telescope; any signal with a drift rate of 0 Hz s-1 is consequently rejected. The second is a spatial filter, motivated by the assumption that extraterrestrial signals should be localized on the sky whereas RFI often is not. Observations of a target are made in a cadence of three “on” pointings interspersed with three “off” pointings in an ABACAD pattern; any signal that appears in the “offs” in addition to the “ons” is likewise rejected. Signals that pass these criteria are promoted from “hits” to “events.”
@@
In this work, we present GLOBULAR (Grouping Low-frequency Observations By Unsupervised Learning After Rejection) clustering, a method that leverages unsupervised learning to reduce false positives in technosignature searches before the spatial filter step by identifying common types of RFI at high spectral resolution and removing them from the data set. In past work, including Mesarcik et al. (2022), signals are not detected a priori, and so noise makes up the clusters while RFI signals are treated as outliers. By contrast, the GLOBULAR clustering pipeline builds off of existing detection algorithms. Once detections have already been made, common RFI types make up the clusters, and unusual RFI signals, as well as any potential extraterrestrial technosignatures, are the outliers.
@@ 1.1 Unsupervised Learning
Our approach implements the HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) algorithm, first proposed by Campello et al. (2013) as an extension of DBSCAN (Ester et al., 1996), and implemented for Python by McInnes et al. (2017). HDBSCAN comes with a number of unique advantages over other clustering algorithms, including outlier detection and the ability to discover the number of clusters without user input. Additionally, unlike DBSCAN, which requires that the user prescribe a distance parameter $\epsilon$, HDBSCAN discovers clusters of different densities in a single run by varying over a preset range of values for $\epsilon$. However, a significant drawback, especially for radio astronomy applications, is that with a complexity $\mathcal{O}(n^{2})$ that can be accelerated only up to $\mathcal{O}(N\log N)$ for certain data sets (McInnes & Healy, 2017), HDBSCAN does not upscale easily to large data volumes. We tackle this problem with a novel batching scheme described in Section 3.1.
@@ 1.2 Data Sample
We also generate 100 synthetic events with setigen (Brzycki et al., 2022) with drift rates between $-3$ and $3$ Hz s-1 and SNRs between 1,000 and 10,000, corresponding to 300 hits since each event appears in three “on” pointings. By attempting to recover these signals, we estimate the algorithm’s true-positive rate. We inject these events into the observation cadence for M 31, found during our toy-model experimentation (Section 4) to be an good example of a typically RFI-dense $L$-band observation, and run a turboSETI FindDoppler search to retrieve the hits. Only 296 of the 300 hits are recovered in this way for analysis in the next stage of the process, not an unexpected result as turboSETI routinely misses a small proportion of signals due to high drift rate, low power, intense RFI in nearby frequency channels, or other reasons. The four missed hits are in two pairs from two different events, so we have a theoretical recovery ceiling of 98 events.
@@ 2 Feature Space
The features used in this work to represent signals are a set of prescriptive statistical parameters that were chosen due to their known relevance to the field. This is in contrast to the “latent space” of features produced by, for example, principal component analysis or the first half of an autoencoder (e.g., Ma et al., 2023), which discover features algorithmically. A benefit to our approach is that the features have clear statistical and/or astrophysical significance to a human observer. While we generally do not know the features’ relative importances — i.e., the degree to which each feature influences the classification of a hit toward one cluster over another — prior to setting them, we can estimate them a posteriori in various ways, including a random forest trained on the clustering results (Section 4.1) and Shapley values (Section 4.2).
Our feature space for this work is 13-dimensional, with 3 of the features calculated natively by the turboSETI implementation and the other 10 calculated afterward. Our feature list is effective, but not exhaustive, and further work may refine the list by introducing new features or pruning relatively unimportant ones as necessary, especially for applications to other telescopes or observing bands. We pre-process our calculated features to ensure that the scales are comparable from one feature to another. Features #1 and 2 undergo quantile transforms, further discussed in their respective sections below. Features #3, 5, 8, 10, 11, and 13 are converted to logarithmic scales. Features #1, 3, 4, 5, 8, 10, 11, 12, and 13 are normalized to unit range, while Feature #2 is normalized to unit variance following its quantile transform to a normal distribution and Feature #9 is normalized to unit maximum with negative values permitted. Features #6 and 7 do not undergo normalization because their ranges are already close to unity. The range of any feature can be adjusted to induce a corresponding effect on the feature’s relative importance; we find, as shown in Sections 4.1 and 4.2, that these ranges are effective for our clustering.
@@
Figure 1: Distributions for the 13 features in our feature space, showing in light blue the distributions before any cluster-based rejection, in medium blue the distributions after 8 epochs of rejection, and in dark blue the retrieved distributions of the injected synthetic signals. Note the logarithmic scale on all $y$-axes. These distributions are from before the pre-processing steps outlined in Section 2, and therefore are not the distributions seen by the HDBSCAN algorithm, which are normalized to ranges of order unity (with the exceptions of Features #6 and 7, which already have ranges of order unity). The gap in observing frequency between 1200 and 1340 MHz is due to a notch filter, the effects of which are removed in pre-processing.
@@
We pre-process the observation frequencies with a quantile transform to a uniform distribution, preserving the signals’ relative locations in frequency space but eliminating those locations’ relative densities. This preserves the relevant information, the apparent spectral proximity of morphologically similar signals to each other, while smoothing out the precise divisions set by the FCC allocations.
#2: Drift rate. The drift rate is calculated by turboSETI and recorded in Hz s-1. A high density of signals near zero drift will occur in virtually every data set due to the time-invariant nature of most RFI. If the density in this region is enormous compared to the wings of the distribution, these signals may be seen as a cluster and the relatively few signals in the wings as outliers, generally an incorrect interpretation. We pre-process the drift rates with a quantile transform to a Gaussian distribution to decrease the likelihood of this interpretation, aiding the algorithm’s identification of low-drift anomalies at the expense of some high-drift sensitivity. This trade-off is desirable because, mainly due to the presence of high-drift transmitters (e.g., Iridium satellites) with a greater drift range than low-drift transmitters, the precise value of a drift rate is less important at high drifts than at low drifts.
@@ 3 Method
Due to the dimensionality of the problem in both feature count (13) and sample size ($\sim 2\times 10^{6}$), the application of HDBSCAN to a large survey of narrowband radio signals is not trivial, requiring certain pre- and post-processing measures to obtain reasonable results in both anomaly detection and RFI classification. A visual summary of these measures is given in Figure 2.
@@
The two dimensions of the embedding contain the information of the input 6-dimensional space as the result of a non-linear mapping, and so have no intuitive physical meaning in either unit or scale. A benefit of t-SNE is that by tuning hyperparameters, the user can place varying degrees of emphasis on local or global structure, seeking a middle ground that emphasizes divisions between groups of clusters and aids matching. Empirically, we find substantial improvements in matching when we employ these dimensionality reduction measures, including improvements in the embedding when it is calculated from a reduced 6-dimensional space instead of the original 13-dimensional space.
@@ 3.1 Batching and Epochal Iteration
In our testing, HDBSCAN lost sensitivity at high data volumes, having a tendency to group an extreme majority ($>$90%) of hits in one cluster. We combat this problem by pre-batching our data and running HDBSCAN on a set of only $\sim$3,000 hits at a time, randomly sampled from the pre-processed ensemble. For our data set, this results in 639 batches. This approach risks splitting up small clusters into pieces too small to be caught by the algorithm, so we iteratively pool the anomalies from each batch into a new data set and run the algorithm again for further reduction.
We deployed our method on the $L$-band subset of the observations used by Choza et al. (2024), a total of 2,186,151 hits (1,917,903 after our excisions and 1,918,199 after our injections; see Section 1.2). The full data set is pre-processed before batching to ensure comparable ranges for all features, and batches are drawn at random from the ensemble with each hit having an equal probability of being drawn. Pre-processing helps ensure that all features are scaled proportionally and that the scaling is not unique on a per-batch basis.
To each batch, we inject an identical set of $100\times 13$ features derived from 100 drifting narrowband signals synthesized with setigen. These signals are distinct from the 100 events injected for recovery; the purpose of these is to “seed” clusters with morphologies we want to keep. Real drifting narrowband signals in the data set are likely to be clustered with the seed signals, enabling the user to more effectively track them through the epochal iteration. Any non-synthetic signals placed in seeded clusters are retained for the next epoch as though they were outliers.
@@ 3.2 Cross-Batch Cluster Matching
Figure 4: Top: 2-dimensional t-SNE embedding of a 10,000-hit sample. The axes are derived from a non-linear mapping and have no intuitive physical meaning. Each hit is passed through t-SNE not with its actual position in our feature space, but with the centroid of the cluster it was found in during the 8 epochs of batching. We run HDBSCAN on the t-SNE embedding to identify which centroids correspond to morphologically similar clusters. HDBSCAN clusters are color-coded by index, with hits in the outlier cluster (indexed $-1$) represented by light gray $\times$’s. Bottom: Histogram showing hit counts for each cluster.
@@
Since distances in a t-SNE embedding are the result of a non-linear mapping and do not necessarily have an intuitive distance-like meaning, post-embedding clusters must be interpreted with caution. Relatively low-level structure in the embedding may in fact be important, yet may be ignored by HDBSCAN in favor of high-level clusters. Conversely, HDBSCAN may find several small clusters where a single supercluster would be more appropriate. There is a trade-off between cluster homogeneity and cluster redundancy. With the embedding in Figure 4, we have chosen to permit some redundancy in the interest of greater homogeneity.

### v2

# Anomaly Detection and Radio-frequency Interference Classification with Unsupervised Learning in Narrowband Radio Technosignature Searches

@@ Abstract
The search for radio technosignatures is an anomaly detection problem: Candidate signals represent needles of interest in the proverbial haystack of radio-frequency interference (RFI). Current search frameworks find an enormity of false-positive signals, especially in large surveys, requiring manual follow-up to a sometimes prohibitive degree. Unsupervised learning provides an algorithmic way to winnow the most anomalous signals from the chaff, as well as group together RFI signals that bear morphological similarities. We present GLOBULAR (Grouping Low-frequency Observations By Unsupervised Learning After Reduction) clustering, a signal processing method that uses HDBSCAN to reduce the false-positive rate and isolate outlier signals for further analysis. When combined with a standard narrowband signal detection and spatial filtering pipeline, such as turboSETI, GLOBULAR clustering offers significant improvements in the false-positive rate over the standard pipeline alone, suggesting dramatic potential for the amelioration of manual follow-up requirements for future large surveys. By removing RFI signals in regions of high spectral occupancy, GLOBULAR clustering may also enable the detection of signals missed by the standard pipeline. We benchmark our method against the Choza et al. turboSETI-only search of 97 nearby galaxies at $L$-band, demonstrating a false-positive hit reduction rate of 93.1% and a false-positive event reduction rate of 99.3%.
@@
software: NumPy (Harris et al., 2020), Matplotlib (Hunter, 2007), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), SciPy (Virtanen et al., 2020), Pandas (McKinney, 2010), h5py (Collette, 2013), Scikit-Learn (Pedregosa et al., 2011), HDBSCAN (McInnes et al., 2017), Blimpy (Price et al., 2019), turboSETI (Enriquez & Price, 2019), setigen (Brzycki et al., 2022), SHAP (Lundberg & Lee, 2017)
@@ 1 Introduction
Since the work of Cocconi & Morrison (1959) and Drake (1961), radio frequencies have comprised the most widely explored domain in the search for extraterrestrial intelligence (SETI) due to their favorably low extinction across cosmic distances and the known efficacy of their use for communication by at least one spacefaring civilization—humanity. Since 2015, the Breakthrough Listen (BL) Initiative (Worden et al., 2017) has used various facilities, including the Robert C. Byrd Green Bank Telescope (GBT), to conduct radio-frequency searches of numerous targets, ranging in scale from the planetary (e.g., Traas et al., 2021; Franz et al., 2022) to the galactic (e.g., Gajjar et al., 2021; Choza et al., 2024), for unambiguously artificial signals, or “technosignatures.” Such signals are generally assumed to be narrowband, with signal widths on the order of 1–10 Hz, for transmission efficiency and since no known astrophysical process naturally produces emission so narrowly peaked. Continuous-wave emission is also typically assumed for feasibility of follow-up observations and while calculating transmitter distributions. Technosignature searches outside the radio have sought other types of signals, such as optical laser line emission (e.g., Zuckerman et al., 2023) and infrared waste heat signatures (e.g., Wright et al., 2014; Suazo et al., 2024). Additionally, some radio technosignature searches have relaxed the narrowband assumption or the continuous-wave assumption, such as, respectively, the searches by Gajjar et al. (2022) for broadband pulsed signals with artificial dispersions and by Suresh et al. (2023) for intermittent signals with artificial periodicities. However, these alternative premises necessitate different search methodologies, and only the narrowband, continuous-wave radio case is treated here.
@@
Past GBT searches (e.g., Enriquez et al., 2017; Price et al., 2020) have primarily used two criteria to reject RFI. The first is that an extraterrestrial narrowband signal must exhibit a drift in frequency over time due to the relative motion between the transmitter and the telescope; any signal with a drift rate of 0 Hz s-1 is consequently rejected. The second is a spatial filter, motivated by the assumption that extraterrestrial signals should be localized on the sky whereas RFI often is not. Observations of a target are made in a cadence of three “on” pointings interspersed with three “off” pointings in an ABACAD pattern, where A designates an on-target scan and B, C, and D designate distinct off-target scans; any signal that appears in the “offs” in addition to the “ons” is likewise rejected. (For a full overview of the BL observation strategy, see Section 2.1 in Lebofsky et al. (2019).) Signals that pass these criteria are promoted from “hits” to “events.”
@@
In this work, we present GLOBULAR (Grouping Low-frequency Observations By Unsupervised Learning After Rejection) clustering, [^1] https://github.com/bjacobell/gbt-hdbscan a method that leverages unsupervised learning to reduce false positives in technosignature searches before the spatial filter step by identifying common types of RFI at high spectral resolution and removing them from the data set. In past work, including Mesarcik et al. (2022), signals are not detected a priori, and so noise makes up the clusters while RFI signals are treated as outliers. By contrast, the GLOBULAR clustering pipeline builds off of existing detection algorithms. Once detections have already been made, common RFI types make up the clusters, and unusual RFI signals, as well as any potential extraterrestrial technosignatures, are the outliers.
@@ 1.1 Unsupervised Learning
Our approach implements the HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) algorithm, first proposed by Campello et al. (2013) as an extension of DBSCAN (Ester et al., 1996), and implemented for Python by McInnes et al. (2017). HDBSCAN comes with a number of unique advantages over other clustering algorithms, including outlier detection and the ability to discover the number of clusters without user input. Additionally, unlike DBSCAN, which requires that the user prescribe a distance parameter $\epsilon$, HDBSCAN discovers clusters of different densities in a single run by varying over a preset range of values for $\epsilon$. However, a significant drawback, especially for radio astronomy applications, is that with a complexity $\mathcal{O}(N^{2})$ that can be accelerated only up to $\mathcal{O}(N\log N)$ for certain data sets (McInnes & Healy, 2017), HDBSCAN does not upscale easily to large data volumes. Previous methods of tackling the runtime problem include an implementation that uses the MapReduce framework across subsamples of the data set with a subsample size determined by single-unit processing power (dos Santos et al., 2019). We employ the batching scheme described in Section 3.1, which uses a subsample size determined by a preliminary analysis of the data, among other differences.
@@ 1.2 Data Sample
Although extragalactic technosignature searches require that any putative transmitters demonstrate truly prodigious power output to be detected, they constrain the existence of such transmitters by sampling trillions of star systems in a single survey (see discussion in Choza et al.). Extragalactic searches may also support unique RFI rejection methods, such as an extension of the scintillation filter used for a Galactic Center search by Brzycki et al. (2024), though these methods were not used by Choza et al. and are not further discussed in this paper.
@@
We also generate 100 synthetic events with setigen (Brzycki et al., 2022) with drift rates between $-3$ and $3$ Hz s-1 and SNRs between 1,000 and 10,000, corresponding to 300 hits since each event appears in three “on” pointings. These signals are designed to resemble the drifting, narrowband, spatially localized signals we aim to discover, modeled after observations of human-made spacecraft (e.g., Voyager 1) and expected to be similar for non-anthropogenic sources. By attempting to recover these signals, we estimate the algorithm’s true-positive rate. We inject these events into the observation cadence for M 31, found during our toy-model experimentation (Section 4) to be an good example of a typically RFI-dense $L$-band observation, and run a turboSETI FindDoppler search to retrieve the hits. Only 296 of the 300 hits are recovered in this way for analysis in the next stage of the process, not an unexpected result as turboSETI routinely misses a small proportion of signals due to high drift rate, low power, intense RFI in nearby frequency channels, or other reasons. The four missed hits are in two pairs from two different events, so we have a theoretical recovery ceiling of 98 events.
@@ 2 Feature Space
The features used in this work to represent signals are a set of prescriptive statistical parameters that were chosen due to their known relevance to the field. This is in contrast to the “latent space” of features produced by, for example, principal component analysis or the first half of an autoencoder (e.g., Ma et al., 2023), which discover features algorithmically. The exact list was developed during preliminary analysis of a small portion of the data set (see Section 4). A benefit to our approach is that the features have clear statistical and/or astrophysical significance to a human observer. While we generally do not know the features’ relative importances — i.e., the degree to which each feature influences the classification of a hit toward one cluster over another — prior to setting them, we can estimate them a posteriori in various ways, including a random forest trained on the clustering results (Section 4.1) and Shapley values (Section 4.2). Our list is intended to be representative of the data rather than tailored specifically to any one algorithm, so while we prioritize HDBSCAN in this work, it may also provide a useful basis for methods such as isolation forests Liu et al. (2008) or unsupervised random forests (e.g., Mantero & Ishwaran, 2021).
Our feature space for this work is 13-dimensional, with 3 of the features calculated natively by the turboSETI implementation and the other 10 calculated afterward. Our feature list is effective, as we show through the feature importance analysis described above and in Sections 4.1 and 4.2, but it is not exhaustive, and further work may refine the list by introducing new features or pruning relatively unimportant ones as necessary, especially for applications to other telescopes or observing bands.
We preprocess our calculated features to ensure that the scales are comparable from one feature to another. Features #1 and 2 undergo quantile transforms, further discussed in their respective sections below. Features #3, 5, 8, 10, 11, and 13 are converted to logarithmic scales. Features #1, 3, 4, 5, 8, 10, 11, 12, and 13 are normalized to unit range, while Feature #2 is normalized to unit variance following its quantile transform to a normal distribution and Feature #9 is normalized to unit maximum with negative values permitted. Features #6 and 7 do not undergo normalization because their ranges are already close to unity. The range of any feature can be adjusted to induce a corresponding effect on the feature’s relative importance; we find, as shown in Sections 4.1 and 4.2, that these ranges are effective for our clustering.
@@
Figure 1: Distributions for the 13 features in our feature space, showing in light blue the distributions before any cluster-based rejection, in medium blue the distributions after 8 epochs of rejection, and in dark blue the retrieved distributions of the injected synthetic signals. Note the logarithmic scale on all $y$-axes. These distributions are from before the preprocessing steps outlined in Section 2, and therefore are not the distributions seen by the HDBSCAN algorithm, which are normalized to ranges of order unity (with the exceptions of Features #6 and 7, which already have ranges of order unity). The gap in observing frequency between 1200 and 1340 MHz is due to a notch filter, the effects of which are removed in preprocessing.
@@
We preprocess the observation frequencies with a quantile transform to a uniform distribution, preserving the signals’ relative locations in frequency space but eliminating those locations’ relative densities. This preserves the relevant information, the apparent spectral proximity of morphologically similar signals to each other, while smoothing out the precise divisions set by the FCC allocations.
#2: Drift rate. The drift rate is calculated by turboSETI and recorded in Hz s-1. A high density of signals near zero drift will occur in virtually every data set due to the time-invariant nature of most RFI. If the density in this region is enormous compared to the wings of the distribution, these signals may be seen as a cluster and the relatively few signals in the wings as outliers, generally an incorrect interpretation. We preprocess the drift rates with a quantile transform to a Gaussian distribution to decrease the likelihood of this interpretation, aiding the algorithm’s identification of low-drift anomalies at the expense of some high-drift sensitivity. This trade-off is desirable because, mainly due to the presence of high-drift transmitters (e.g., Iridium satellites) with a greater drift range than low-drift transmitters, the precise value of a drift rate is less important at high drifts than at low drifts.
@@
Limitations of the bimodality coefficient in statistical inference are well documented, including that its probability density function is unknown and so statistical significance tests cannot be applied (e.g., Knapp, 2007; Pfister et al., 2013). The dip test of Hartigan & Hartigan (1985) may provide a viable alternative for inference analysis using these data.
@@ 3 Method
Due to the dimensionality of the problem in both feature count (13) and sample size ($\sim 2\times 10^{6}$), the application of HDBSCAN to a large survey of narrowband radio signals is not trivial, requiring certain pre- and postprocessing measures to obtain reasonable results in both anomaly detection and RFI classification. A visual summary of these measures is given in Figure 2.
@@
The two dimensions of the embedding contain the information of the input 6-dimensional space as the result of a non-linear mapping, and so have no intuitive physical meaning in either unit or scale. A benefit of t-SNE is that by tuning hyperparameters, the user can place varying degrees of emphasis on local or global structure, seeking a middle ground that emphasizes divisions between groups of clusters and aids matching. However, the ideal tuning is subjective, and, combined with the difficulty of interpreting apparent structures in t-SNE embeddings (Wattenberg et al., 2016), can easily lead to faulty results. We emphasize that visually spot-checking the clusters found by HDBSCAN at this stage is valuable for verifying that the output is sensible. We also note that empirically, we find substantial improvements in matching when we employ the dimensionality reduction measures discussed above, including improvements in the embedding when it is calculated from a reduced 6-dimensional space instead of the original 13-dimensional space.
@@ 3.1 Batching and Epochal Iteration
In our testing, HDBSCAN lost sensitivity at high data volumes, having a tendency to group an extreme majority ($>$90%) of hits in one cluster. We combat this problem by pre-batching our data and running HDBSCAN on a set of only $\sim$3,000 hits at a time, randomly sampled from the preprocessed ensemble. For our data set, this results in 639 batches. This approach risks splitting up small clusters into pieces too small to be caught by the algorithm, so we iteratively pool the anomalies from each batch into a new data set and run the algorithm again for further reduction.
We deployed our method on the $L$-band subset of the observations used by Choza et al. (2024), a total of 2,186,151 hits (1,917,903 after our excisions and 1,918,199 after our injections; see Section 1.2). The full data set is preprocessed before batching to ensure comparable ranges for all features, and batches are drawn at random from the ensemble with each hit having an equal probability of being drawn. Preprocessing helps ensure that all features are scaled proportionally and that the scaling is not unique on a per-batch basis.
To each batch, we inject an identical set of $100\times 13$ features derived from 100 drifting narrowband signals synthesized with setigen. These signals are generated in the same way as, but are otherwise distinct from, the 100 events injected for recovery; the purpose of these is to “seed” clusters with morphologies we want to keep. Real drifting narrowband signals in the data set are likely to be clustered with the seed signals, enabling the user to more effectively track them through the epochal iteration. Any non-synthetic signals placed in seeded clusters are retained for the next epoch as though they were outliers.
@@ 3.2 Cross-Batch Cluster Matching
Figure 4: Top: 2-dimensional t-SNE embedding of a 10,000-hit sample. The axes are derived from a non-linear mapping and have no intuitive physical meaning. Each hit is passed through t-SNE not with its actual position in our feature space, but with the centroid of the cluster it was found in during the 8 epochs of batching. We run HDBSCAN on the t-SNE embedding to identify which centroids correspond to morphologically similar clusters. HDBSCAN clusters are color-coded by index, with hits in the outlier cluster (indexed $-1$) represented by light gray $\times$’s. The 59 clusters shown here are an estimate, not a 1-to-1 identification, of common RFI sources in the environment. Bottom: Histogram showing hit counts for each cluster.
@@
Since distances in a t-SNE embedding are the result of a non-linear mapping and do not necessarily have an intuitive distance-like meaning, apparent clusters in the embedding must be interpreted with caution. Relatively low-level structure in the embedding may in fact be important, yet may be ignored by HDBSCAN in favor of high-level clusters. Conversely, HDBSCAN may find several small clusters where a single supercluster would be more appropriate. There is a trade-off between cluster homogeneity and cluster redundancy. With the embedding in Figure 4, we have chosen to permit some redundancy in the interest of greater homogeneity.